# Probabilistic estimates of the diameters of the Rubik’s Cube groups

### v6

@@ Abstract
The diameter of the Cayley graph of the Rubik’s Cube group is the fewest number of turns needed to solve the Cube from the hardest initial configuration. For the 2$\times$2$\times$2 Cube, the diameter is 11 in the half-turn metric, 14 in the quarter-turn metric, 19 in the semi-quarter-turn metric, and 10 in the bi-quarter-turn metric. For the 3$\times$3$\times$3 Cube, the diameter was determined by Rokicki et al.  to be 20 in the half-turn metric and 26 in the quarter-turn metric. This study shows that a modified version of the coupon collector’s problem in probabilistic theory can predict the diameters correctly for both 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes insofar as the quarter-turn metric is adopted. In the half-turn metric, the diameters are overestimated by one and two, respectively, for the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes, whereas for the 2$\times$2$\times$2 Cube in the semi-quarter-turn and bi-quarter-turn metrics, they are overestimated by two and underestimated by one, respectively. Invoking the same probabilistic logic, the diameters of the 4$\times$4$\times$4 and 5$\times$5$\times$5 Cubes are predicted to be 48 (41) and 68 (58) in the quarter-turn (half-turn) metric, whose precise determinations are far beyond reach of classical supercomputing. The probabilistically estimated diameter is shown to obey the approximate formula of $\ln N/\ln r+\ln N/r$, where $N$ is the number of configurations and $r$ is the branching ratio.
@@ III Probabilistic estimate
We shall instead estimate this number by simulating the above brute-force algorithm under the assumption that all generated configurations are random and independent of the previous “seed” configurations (this non-Markovian assumption is clearly not obeyed in reality and is a major approximation). We can then view this problem as a modified version of the coupon collector’s problem Gumbel (1941); Dawkins (1991); Flajolet et al. (1992); Blom et al. (1994) in probabilistic theory (see Appendix A). See pages 26 and 34 of Singmaster’s book Singmaster (1981), where the author develops a similar argument. It is also related to mathematical models of epidemics Hill (2020) and network theory Newman (2008).
@@ IV The 2$\times$2$\times$2 Cube
One of the eight corner cubies (say, the front-top-left one) of the 2$\times$2$\times$2 Cube is held fixed as the orientation anchor of the whole Cube (here, a “cubie” stands for a 1$\times$1$\times$1 constituent cube). The remaining seven cubies’ positions can be permuted in $7!$ ways and their orientations can take $3^{7}$ ways, except the last (seventh) cubie’s orientation is fixed by the first six (it is said Hofstadter (1981) that the “sanity” of the seven cubies is maintained by only one of three orbits Joyner (2008)). Therefore, the total number of configurations $N$ (or the order of the 2$\times$2$\times$2 Rubik’s Cube group) is
@@ IV.1 Half-turn metric
With the front-top-left cubie held fixed, only the right ($R$), downward ($D$) and back ($B$) layers can be rotated. In the half-turn metric, the Cube can undergo one of the three (3) clockwise 90 turns ($R$, $D$, $B$), three (3) counterclockwise 90 turns ($R^{-1}$, $D^{-1}$, $B^{-1}$), and three (3) 180 turns ($R^{2}$, $D^{2}$, $B^{2}$) in the Singmaster notation Singmaster (1981).
@@ V.3 Square-turn metric
The first factor of $4!$ is the number of ways in which the urf corner cubie Singmaster (1981) and its three diagonal ones (ulf, drb, dlb) can interchange their positions by the $R^{2}$, $D^{2}$, and $B^{2}$ turns (their orientations are fixed by their positions). The second factor of $4$ is the number of positions the ulf cubie can reach by the $L^{2}$, $U^{2}$, and $F^{2}$ turns. The remaining three corner cubies’ positions and orientations are uniquely determined by the first five. The quotient $4!4!4!/2$ is the number of ways three groups of 4 edge cubies (sharing a common center slice) can be positioned with their orientations fixed. Since the number of correctly oriented edge cubies is always even, there are two orbits and the number must be divided by two.
@@ VIII.2 Size dependence
In accompanying article Hirata (2024b), a lower bound for the diameter of symmetric Cayley graphs of some Rubik’s Cube groups is obtained on the basis of graph-theoretical arguments, and compared with lower and upper bounds for the diameter of random regular graphs of Bollobás and de la Vega Bollobás and de la Vega (1982); Bollobás (2001), who combined graph-theoretical and probabilistic arguments.

### v7

@@ Abstract
The diameter of the Cayley graph of the Rubik’s Cube group is the fewest number of turns needed to solve the Cube from the hardest initial configuration. For the 2$\times$2$\times$2 Cube, the diameter is 11 in the half-turn metric, 14 in the quarter-turn metric, 19 in the semi-quarter-turn metric, and 10 in the bi-quarter-turn metric. For the 3$\times$3$\times$3 Cube, the diameter was determined by Rokicki et al.  to be 20 in the half-turn metric and 26 in the quarter-turn metric. This study shows that a modified version of the coupon collector’s problem in probability theory can predict the diameters correctly for both 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes insofar as the quarter-turn metric is adopted. In the half-turn metric, the diameters are overestimated by one and two, respectively, for the 2$\times$2$\times$2 and 3$\times$3$\times$3 Cubes, whereas for the 2$\times$2$\times$2 Cube in the semi-quarter-turn and bi-quarter-turn metrics, they are overestimated by two and underestimated by one, respectively. Invoking the same probabilistic logic, the diameters of the 4$\times$4$\times$4 and 5$\times$5$\times$5 Cubes are predicted to be 48 (41) and 68 (58) in the quarter-turn (half-turn) metric, whose precise determinations are far beyond reach of classical supercomputing. The probabilistically estimated diameter is shown to obey the approximate formula of $\ln N/\ln r+\ln N/r$, where $N$ is the number of configurations and $r$ is the branching ratio.
@@ III Probabilistic estimate
We shall instead estimate this number by simulating the above brute-force algorithm under the assumption that all generated configurations are random and independent of the previous “seed” configurations (this non-Markovian assumption is clearly not obeyed in reality and is a major approximation). We can then view this problem as a modified version of the coupon collector’s problem Gumbel (1941); Dawkins (1991); Flajolet et al. (1992); Blom et al. (1994) in probability theory (see Appendix A). See pages 26 and 34 of Singmaster’s book Singmaster (1981), where the author develops a similar argument. It is also related to mathematical models of epidemics Hill (2020) and network theory Newman (2008).
@@ IV The 2$\times$2$\times$2 Cube
One of the eight corner cubies (say, the front-up-left one) of the 2$\times$2$\times$2 Cube is held fixed as the orientation anchor of the whole Cube (here, a “cubie” stands for a 1$\times$1$\times$1 constituent cube). The remaining seven cubies’ positions can be permuted in $7!$ ways and their orientations can take $3^{7}$ ways, except the last (seventh) cubie’s orientation is fixed by the first six (it is said Hofstadter (1981) that the “sanity” of the seven cubies is maintained by only one of three orbits Joyner (2008)). Therefore, the total number of configurations $N$ (or the order of the 2$\times$2$\times$2 Rubik’s Cube group) is
@@ IV.1 Half-turn metric
With the front-up-left cubie held fixed, only the right ($R$), downward ($D$) and back ($B$) layers can be rotated. In the half-turn metric, the Cube can undergo one of the three (3) clockwise 90 turns ($R$, $D$, $B$), three (3) counterclockwise 90 turns ($R^{-1}$, $D^{-1}$, $B^{-1}$), and three (3) 180 turns ($R^{2}$, $D^{2}$, $B^{2}$) in the Singmaster notation Singmaster (1981).
@@ V.3 Square-turn metric
The first factor of $4!$ is the number of ways in which the front-up-right (fur) corner cubie Singmaster (1981) and its three diagonal ones (fdl, bdr, bul) can interchange their positions by the $R^{2}$, $D^{2}$, and $B^{2}$ turns (their orientations are fixed by their positions). The second factor of $4$ is the number of positions the front-up-left ful cubie can reach by the $L^{2}$, $U^{2}$, and $F^{2}$ turns. The remaining three corner cubies’ positions and orientations are uniquely determined by the first five. The quotient $4!4!4!/2$ is the number of ways three groups of 4 edge cubies (sharing a common center slice) can be positioned with their orientations fixed. Since the number of correctly oriented edge cubies is always even, there are two orbits and the number must be divided by two.
@@ VIII.2 Size dependence
In the following article Hirata (2024b), a lower bound for the diameter of symmetric Cayley graphs of some Rubik’s Cube groups is obtained on the basis of graph-theoretical arguments, and compared with lower and upper bounds for the diameter of random regular graphs of Bollobás and de la Vega Bollobás and de la Vega (1982); Bollobás (2001), who combined graph-theoretical and probabilistic arguments.